Find the dimensions of the right circular cone of the greatest lateral area that can be inscribed in a sphere of radius α


1
Expert's answer
2021-12-16T17:08:29-0500

Solution:


Given, a cone inscribed in a sphere of radius a.

Let the radius, height and slant height of cone be "r,h,l" respectively.

"OB=OA=a,BC=r,AC=h,OC=h-a,AB=l"

Using pythagoras theorem in triangle BOC,

"BC^2+CO^2=OB^2
\\\Rightarrow r^2+(h-a)^2=a^2
\\\Rightarrow r^2+h^2+a^2-2ha=a^2
\\\Rightarrow r^2+h^2=2ha \ ...(i)"

Using pythagoras theorem in triangle BAC,

"BC^2+CA^2=AB^2
\\\Rightarrow r^2+h^2=l^2\ ...(ii)"

From (i) and (ii)

"2ha=l^2"

Now, lateral area of cone, "L=\pi rl"

"L^2=\pi^2 r^2 l^2
\\=\pi^2 (l^2-h^2)l^2\ [From\ (ii)] 
\\=\pi^2(2ha-h^2)(2ha)
\\=\pi^2(4h^2a^2-2h^3a)"

On differentiating w.r.t. h

"2L\dfrac{dL}{dh}=\pi^2(8ha^2-6h^2a) ...(iii)
\\ \Rightarrow\dfrac{dL}{dh}=\dfrac{\pi^2 (2ha\pi^2 (4a-3h))}{2L}"

Now, put "\dfrac{dL}{dh}=0"

"\\ \Rightarrow\dfrac{\pi^2 (2ha\pi^2 (4a-3h))}{2L}=0
\\ \Rightarrow 4a-3h=0
\\ \Rightarrow h=\dfrac{4a}3\ ...(iv)"

Again, differentiating (iii) w.r.t h,

"2L\dfrac{d^2L}{dh^2}+2(\dfrac{dL}{dh})^2=\pi^2(8a^2-12ha)
\\=\pi^2(8a^2-12a(\dfrac{4a}3))\ [Using\ (iv)]
\\=-8\pi^2a^2<0
\\\Rightarrow\ maxima"

Put (iv) in (i)

"\\r^2+(\dfrac{4a}3)^2=2(\dfrac{4a}3)a 
\\\Rightarrow r^2=\dfrac{8a^2}3-\dfrac{16a^2}9
\\\Rightarrow r^2=\dfrac{8a^2}9
\\\Rightarrow r=\dfrac{2\sqrt2a}3"

Thus, the dimensions of cone are "r=\dfrac{2\sqrt2a}3;h=\dfrac{4a}3"